A 100-point test contains a total of 30 questions. The multiple choice questions are worth 3 points each and the short response questions are worth 5 points each.
Write and solve a system of equations to find the number of MC questions (x) and the number of short response questions (y).

Answer by mananth(16946)

1 x + 1 y = 30


3 x + 5 y = 100 .............2
Eliminate y
multiply (1)by -5
Multiply (2) by 1
-5 x -5 y = -150
3 x 5 y = 100
Add the two equations
-2 x = -50
/ -2
x = 25
plug value of x in (1)
1 x + 1 y = 30
25 + 1 y = 30
1 y = 5
y = 5
Ans x = 25
y = 5
25 Number of multiple choice questions
5 Number of short answer questions